OpenStudy (anonymous):

Auditions for a new, upcoming band in Chicago were held with 90 men and women trying out for the lead singer position. If twice the number of men less three times the number of women is equal to 30, how many women auditioned? 50 40 60 30

OpenStudy (anonymous):

I would be dead without your guyzzz help . <3

OpenStudy (swissgirl):

ok let women = x There x amount of women and let men = (90 - x) since there are 90 ppl in total

OpenStudy (alexwee123):

make a system of equations! let x=#of men and y=#of women so... x+y=90 2x-3y=30

OpenStudy (swissgirl):

ohh yaaa u can solve it that way

OpenStudy (anonymous):

whats the next step?

OpenStudy (alexwee123):

so lets use substitution to solve this make x=90-y and sub that in for x into 2x-3y=30 so... 2(90-y)-3y=30 180-2y-3y=30 -5y=-150 y=30 so 30 women auditioned

OpenStudy (alexwee123):

and if you also need to find out the #of men who auditioned then just simply sub in 30 for y into x+y=90 to get your answer :)